Ariel dropped a golf ball from her second story window. The ball starts from rest and hits the sidewalk 3.5 s later with a velocity of 15.0 m/s. Find the average acceleration of the golf ball.

Bold

Italic

Underline





1
Expert's answer
2022-03-16T18:28:16-0400

The average acceleration of the golf ball

"a=\frac{\Delta v}{\Delta t}=\frac{15-0}{3.5}=4.3\:\rm m/s^2"